# Cautionary tales on heating-rate prescriptions in kilonovae

### v1

@@ Abstract
A major ingredient for kilonova lightcurves is the radioactive heating rate and its dependence on the electron fraction and velocity of the ejecta and, in principle, on the nuclear mass formula. Heating-rate formulae commonly used as the basis for kilonova models are, strictly speaking, incorrect for electron fractions other than $Y_{e}=0.04$. Here, we introduce new semi-analytical models for kilonovae with better heating rate prescriptions valid for the full parameter space of kilonova velocities and electron fractions. This new prescription produces, on average, dimmer kilonovae at peak that decay more slowly as compared to previously used prescriptions for otherwise identical kilonova physics. We show the dangers of using inappropriate heating rate estimates by simulating realistic observations and inferring the kilonova parameters via a misspecified heating-rate prescription. While providing great fits to the photometry, an incorrect heating-rate prescription fails to recover the input ejecta masses at greater than $5\sigma$. This bias from an incorrect prescription has disastrous consequences for interpreting kilonovae, their use as additional components in gamma-ray burst afterglows, and understanding their role in cosmic chemical evolution or for multi-messenger constraints on the nuclear equation of state. Given the true heating-rate is uncertain, we estimate there is a $\approx 5$ and $\approx 10\%$ systematic uncertainty in the measured ejecta masses and velocities, respectively. For lanthanide-rich ejecta, the systematic uncertainty could be as high as $\approx 50\%$. This systematic uncertainty limits the precision of any measurements that rely on accurate estimates of kilonova ejecta properties.
@@ 1 Introduction
The electromagnetic observations of AT2017gfo (Smartt et al., 2017; Villar et al., 2017; Cowperthwaite et al., 2017; Tanvir et al., 2017; Kasen et al., 2017; Evans et al., 2017), the kilonova that accompanied the first gravitational-wave observation of a binary neutron star merger (Abbott et al., 2017a, b), confirmed that binary neutron star mergers are sites of rapid neutron capture (r-process) nucleosynthesis as theoretically expected Lattimer & Schramm (1974); Symbalisty & Schramm (1982); Eichler et al. (1989); Rosswog et al. (1999); Freiburghaus et al. (1999). Detailed spectroscopic observations and modelling would later suggest the production of specific r-process elements such as Strontium (Watson et al., 2019). The combination of electromagnetic and gravitational-wave observations of GW170817 would go on to provide many constraints on significant questions in physics and astrophysics, such as a standard-siren measurement of the Hubble constant, independent of the cosmic distance ladder (Abbott et al., 2017c), constraints on the nuclear equation of state (Margalit & Metzger, 2017; Coughlin et al., 2019; Nicholl et al., 2021), and the confirmation that gravitational waves propagate to an enormous accuracy at the speed of light  (Abbott et al., 2017d).
@@
Figure 1: Set of lightcurves for different Vera Rubin Observatory filters with a two-component kilonova model for different ejecta velocities and electron fractions (other parameters are identical). For all three models we fix the ejecta masses of each component to $0.02$ and $0.04~{}M_{\odot}$, and floor temperatures to $2500$ and $1000$ K, respectively to isolate the impact of the heating rate. The blue and green curves show two component ejecta’s with electron fractions of $0.4$, $0.23$, and $0.25$ and $0.1$, while the red curves show both components for an electron fraction of $0.4$. For the ejecta velocities, we fix the first component velocity to $0.25~{}c$ for all curves, while the second component velocities are $0.1$, $0.15$, and $0.2~{}c$, for the blue, green, and red curves, respectively. The old heating-rate prescription (Korobkin et al., 2012) is shown as dashed curves while the solid curves show the updated prescription (Rosswog & Korobkin, 2024). The physics behind the two models is the same apart from the heating-rate prescription.
@@
The detection and interpretation of kilonovae light curves are heavily reliant on available kilonova models. Specifically, such models form a key aspect for designing surveys, understanding depth requirements for follow-up, and ultimately characterising and inferring the properties of candidate events. However, kilonova light curve models are subject to a number of different ingredients, each with their own underlying uncertainty. One of these critical modelling ingredients is the nuclear heating rate, i.e., the energy produced per second by one gram of r-process material, which is ultimately subject to the decay channels of r-process nuclei (see e.g., Cowan et al., 2021, for an extensive review on various aspects of rapid neutron capture nucleosynthesis).
@@
The temporal decay of a kilonova lightcurve has significant diagnostic power and may provide clues to the ejecta composition. If a broad distribution of nuclei is available and all are decaying exponentially, one expects a power-law dependence of the heating rate as a function of time (Way & Wigner, 1948; Metzger et al., 2010; Hotokezaka et al., 2017). If, instead, only a few isotopes dominate the heating rate, one expects distinctive “bumps” in the lightcurve that can provide constraints on the half-lives of the decaying species (Grossman et al., 2014; Martin et al., 2015; Lippuner & Roberts, 2015; Rosswog et al., 2018; Wanajo, 2018; Wu et al., 2017; Kasliwal et al., 2022).
@@
The specific interactions that take place and their decay products determine how much energy is deposited in the expanding ejecta cloud and the overall heating rate. Any neutrino products escape the ejecta and do not contribute to the overall heating rate, while the heating-rate efficiency of $\gamma-$rays decays very rapidly and becomes negligible after tens of days (Hotokezaka et al., 2016). On the other hand, charged particles, such as $\alpha$-particles and fission products, keep depositing energy into the ejecta even at late times, and this makes kilonova lightcurves sensitive to the contribution of those decay channels (Barnes et al., 2016; Rosswog et al., 2017; Kasen & Barnes, 2019; Wu et al., 2019; Zhu et al., 2018; Vassh et al., 2019; Giuliani et al., 2020).
@@
Understanding what interactions take place and what their decay products are is non-trivial. To complicate matters further, the interactions themselves are sensitive to the overall composition of the kilonova ejecta, which itself is linked in a multitude of ways to the progenitor binary neutron star system, equation of state, and the fate of the remnant (e.g., Rosswog et al., 2018; Sarin & Lasky, 2021). For relatively high $Y_{e}$ values, $\beta$-decays dominate, and only the resulting electrons and photons are relevant for heating the ejecta; again, the neutrinos escape essentially without any interaction. In low $Y_{e}$ ejecta, actinides can be produced which, via $\alpha$-decays and fission, can deposit large amounts of energy, altering the heating rate by up to an order of magnitude depending on the underlying nuclear mass model (Barnes et al., 2016; Rosswog et al., 2017; Wu et al., 2019).
@@
In principle, kilonova models should be calculated self-consistently, such that feedback from the energy released and how this affects the expansion of the ejecta and nucleosynthesis/heating rate is incorporated. Recent work suggests that failure to include this feedback can compromise the abundance evolution (Magistrelli et al., 2024). Moreover, kilonova models should keep track of the exact decay channels to determine both the “naked” heating rate (i.e., the raw output) and the thermalization efficiency (the fraction that is coupled to the ejecta). This is difficult in practice since such models would need to be coupled to nuclear reaction networks. Instead, it is common to construct separate expressions for the naked heating rate, $\dot{q}$, and the thermalization efficiency, $\epsilon_{\rm th}$. The effective heating rate is then $\dot{q}_{\rm eff}=\epsilon_{\rm th}\;\dot{q}$.
@@
$\dot{q}=q_{0}\left(\frac{1}{2}-\frac{1}{\pi}{\rm arctan}\frac{t-t_{0}}{\sigma}\right)^{\alpha},$ (1)
with $q_{0}=2\times 10^{18}$ erg/(g s), $t_{0}=1.3$ s, $\sigma=0.11$ s, $\alpha=1.3$ (Metzger et al., 2010), and the thermalization efficiency, $\epsilon_{\rm th}$ computed following Barnes et al. (2016) or Wollaeger et al. (2018). This specific heating-rate prescription was obtained as a fit to the heating rate from a broad set of Newtonian+GW-backreaction hydrodynamics simulations with the stiff equation of state of Shen et al. (1998a, b) and a multi-flavour leakage scheme (Rosswog & Liebendörfer, 2003). This prescription has served as the basis for a large variety of kilonova models, including the majority of model fits to the lightcurve of AT2017gfo, GRB211211A, and GRB230307A (Villar et al., 2017; Nicholl et al., 2021; Rastinejad et al., 2022; Levan et al., 2024), and detailed radiative transfer simulations (Bulla, 2019).
Figure 2: $68\%$ credible interval from $5000$ simulated kilonovae with fixed ejecta masses and temperature floors and varying velocities and opacities for old prescription in blue and updated prescription in green.
@@
Full general-relativistic simulations with more sophisticated neutrino treatments, produce, as expected, a broader range of electron fractions (Perego et al., 2019; Shibata & Hotokezaka, 2019), which requires significantly different heating-rate prescriptions (Rosswog & Korobkin, 2024). The fit in the latter reference is based on a library of parametrized trajectories where velocity and electron fraction were systematically varied. The actual nuclear network calculations were performed with WinNet (Winteler, 2012; Winteler et al., 2012; Reichert et al., 2023) which is based on the BasNet network (Thielemann et al., 2011). The functional dependence of the fit is more complicated than the above prescription, with multiple additional terms and written as,
@@
where the coefficients no longer have fixed values, but depend on the velocity and electron fraction of the ejecta. We refer to the appendix of Rosswog & Korobkin (2024) for the explicit coefficient expressions. The role of this updated heating-rate prescription has been explored in detailed radiative transfer simulations of kilonovae (Bulla, 2023), or used for predictions for surveys like LSST using the Vera Rubin Observatory (Setzer et al., 2023). However, it is not explored in detail for semi-analytical models. Moreover, it is unclear what the consequences of using the previous heating-rate prescription in fitting light curves are and what the systematic uncertainty for inferences of kilonova ejecta parameters due to uncertainties in the heating rate truly is.
In this Letter, we explore the impact of the updated heating rate prescription of Rosswog & Korobkin (2024). In particular, in Sec. 2, we briefly describe a semi-analytical kilonova model implemented in open-source software, Redback (Sarin et al., 2023), highlighting the differences in individual light curves and a population due to the heating-rate prescription alone. We follow this by showcasing the dangers of using the previous heating-rate prescription to fit realistic, simulated kilonovae observations in Sec. 3. We then discuss the implications of uncertainty in the heating-rate prescriptions and conclude in Secs. 4 and 5, respectively.
@@ 2 Modelling and impact of heating-rate
Early modelling of the photometry of AT2017gfo made it abundantly clear that one single ejecta composition was inadequate for explaining the observations, with a clear evolution from a bright blue to a red source (Cowperthwaite et al., 2017; Kasen et al., 2017). This led to the development of multi-component models, where each distinct ejecta component is characterized by a single opacity (Kasen et al., 2017). For the specific case of AT2017gfo, a two-component model with one lanthanide-poor, blue component and another lanthanide-rich red component proved sufficient to provide good fits to the observations (Villar et al., 2017), with extensions to include the effects of other emission mechanisms such as shock-cooling (Arcavi et al., 2017), effects of the hypermassive neutron star wind (Metzger et al., 2018), or interaction with a jet (Kasliwal et al., 2017).
Although these extensions are relevant, the basic ingredients of the majority of semi-analytical models for kilonovae are broadly similar. We first have the heating-rate prescription, which, for the majority of models follows Eq. 1. Only some fraction of this input energy is available to thermalize this ejecta, which is often taken from Barnes et al. (2016),
@@
Here, a, b, c are functions of the ejecta velocity and mass, with the overall function built as a fit to simulations (Barnes et al., 2016). The overall bolometric luminosity is then estimated following the Arnett model for supernovae (Arnett, 1982), with a temperature and wavelength independent gray opacity, which is estimated from the electron fraction, $Y_{e}$ following Tanaka & Hotokezaka (2013). As most comparisons to data are done through observations in specific filters, an assumption for the spectral energy distribution is also required, with the typical choice being a blackbody with a floor temperature such that the photospheric radius starts receding and the temperature remains constant once the floor temperature is reached.
@@
As we described in the introduction, the heating-rate prescription used by these models is only valid for an electron fraction $Y_{e}=0.04$, i.e., (following the conversion from Tanaka & Hotokezaka (2013)), a grey opacity of $36\,\mathrm{g^{-1}~{}cm^{-2}}$, in stark contrast to the opacities fixed for different components when fitting AT2017gfo, and other kilonova candidates. For example, the two-component fit for AT2017gfo in Villar et al. (2017) had gray opacities of $0.5\,\mathrm{g^{-1}~{}cm^{-2}}$ and $10\,\mathrm{g^{-1}~{}cm^{-2}}$. A different heating-rate prescription has a dramatic impact on the inferred properties of the ejecta (we explore this in more detail in Sec. 3). However, first we investigate the impact of the updated heating-rate prescription on the kilonova lightcurves themselves.
@@
In Fig. 1, we show two representative two component kilonova model magnitudes at $40\,\mathrm{Mpc}$. Here the difference is only due to the heating-rate prescription, with the old prescription (Korobkin et al., 2012) shown as dashed curves while the solid curve is the updated prescription (Rosswog & Korobkin, 2024). For all three models we fix the ejecta masses of each component to $0.02$ and $0.04~{}M_{\odot}$, and floor temperatures to $2500$ and $1000$ K, respectively to isolate the impact of the heating rate. The blue and green curves show two component ejecta’s with electron fractions of $0.4$, $0.23$, and $0.25$ and $0.1$, while the red curves show both components for an electron fraction of $0.4$. For the ejecta velocities, we fix the first component velocity to $0.25~{}c$ for all curves, while the second component velocities are $0.1$, $0.15$, and $0.2~{}c$, for the blue, green, and red curves, respectively. These curves demonstrate the significant impact of the heating-rate prescription, not only in changing the magnitude at peak by up to $0.5$ mag, but also the temporal evolution of the lightcurve. In general, the previous prescription produces brighter kilonovae at peak, but the late-time evolution is generally faster, although the parameters for the red curve indicate the new prescription is brighter, with a significant deviation post $\approx 5\,\mathrm{d}$, entirely due to the differences in input heating rate.
@@
Although the above detailed lightcurves highlight specific cases where the updated heating-rate prescription leads to a markedly different lightcurve. It is also instructive, to look at a broader set of lightcurves. To understand the impact across a broad sample, we create $5000$ lightcurves for the same parameters randomly drawn from uniform priors on opacity between $1$ and $30\,\mathrm{g^{-1}~{}cm^{-2}}$. Truncated normal priors for velocity with the first component with $\mu=0.15$, $\sigma=0.1$c, and the second component with $\mu=0.1$, $\sigma=0.05$c. For both components we fix the minimum and maximum velocities to $0.05$ and $0.35$c, respectively. We also fix the ejecta masses and temperature floors to $0.01$, $0.03~{}M_{\odot}$ and $3500$ and $1500$ K, again to isolate the impact of the heating-rate prescription. In reality, the properties of the ejecta, such as the masses, opacities, velocities are all correlated and link to the properties of the binary, e.g., through relations calibrated to numerical relativity simulations (Coughlin et al., 2019; Setzer et al., 2023). However, these relations are largely uncertain and can at times make nonsensical predictions outside the regime of the small subset of numerical simulations they are based on (Henkel et al., 2023), something we wish to avoid.
@@
In Fig. 2, we show the $68\%$ credible interval from the sample of $5000$ lightcurves, with the old prescription in blue and the updated heating-rate prescription in green. These lightcurve distributions highlight the broad impact of the updated heating-rate prescription, with the new prescription in general result in a dimmer kilonova at peak but one that declines more slowly. Another feature these lightcurve distributions highlight is the significantly tighter shape of the previous prescription (blue) i.e., the evolution of the lightcurve with the previous prescription is largely similar regardless of the opacity or ejecta velocity. This is in stark contrast to the updated prescription, where the lightcurve evolution depends sensitively on the opacity and ejecta velocity, a direct consequence of their impact on the input heating-rate.

### v2

@@ Abstract
A major ingredient for kilonova lightcurves is the radioactive heating rate and its dependence on the electron fraction and velocity of the ejecta and, in principle, on the nuclear mass formula. Heating-rate formulae commonly used as the basis for kilonova models previously employed in the literature produce substantially different outputs for high electron fractions ($Y_{e}\gtrsim 0.3$) and at late times ($t\gtrsim 1\,\mathrm{d}$) compared to newer prescriptions. Here, we employ standard semi-analytical models for kilonovae with better heating rate prescriptions valid for the full parameter space of kilonova velocities and electron fractions to explore the impact of the heating rate on kilonova lightcurves. We show the dangers of using inappropriate heating rate estimates by simulating realistic observations and inferring the kilonova parameters via a misspecified heating-rate prescription. While providing great fits to the photometry, an incorrect heating-rate prescription fails to recover the input ejecta masses with a bias significantly larger than the typical statistical uncertainty. This bias from an incorrect prescription has significant consequences for interpreting kilonovae, their use as additional components in gamma-ray burst afterglows, and understanding their role in cosmic chemical evolution or for multi-messenger constraints on the nuclear equation of state. We showcase a framework and tool to better determine the impact of different modelling assumptions and uncertainties on inferences into kilonova properties.
@@ 1 Introduction
The electromagnetic observations of AT2017gfo (Smartt et al., 2017; Villar et al., 2017; Cowperthwaite et al., 2017; Tanvir et al., 2017; Kasen et al., 2017; Evans et al., 2017), the kilonova that accompanied the first gravitational-wave observation of a binary neutron star merger (Abbott et al., 2017a, b), confirmed that binary neutron star mergers are sites of rapid neutron capture (r-process) nucleosynthesis as theoretically expected (Lattimer & Schramm, 1974; Symbalisty & Schramm, 1982; Eichler et al., 1989; Rosswog et al., 1999; Freiburghaus et al., 1999). Detailed spectroscopic observations and modelling would later suggest the production of specific r-process elements such as Strontium (Watson et al., 2019). The combination of electromagnetic and gravitational-wave observations of GW170817 would go on to provide many constraints on significant questions in physics and astrophysics, such as a standard-siren measurement of the Hubble constant, independent of the cosmic distance ladder (Abbott et al., 2017c), constraints on the nuclear equation of state (Margalit & Metzger, 2017; Coughlin et al., 2019; Nicholl et al., 2021), and the confirmation that gravitational waves propagate to an enormous accuracy at the speed of light  (Abbott et al., 2017d).
@@
Figure 1: Set of lightcurves for different Vera Rubin Observatory filters with a two-component kilonova model for different ejecta velocities and electron fractions (other parameters are identical). For all three models we fix the ejecta masses of each component to $0.02$ and $0.04~{}M_{\odot}$, and floor temperatures to $1500$ and $500$ K, respectively to isolate the impact of the heating rate. The blue and green curves show two component ejecta’s with electron fractions of $0.35$, $0.05$, and $0.25$ and $0.1$, while the red curves show both components for an electron fraction of $0.4$. For the ejecta velocities, we fix the first component velocity to $0.05~{}c$ for the blue and red curves and $0.25~{}c$ for the green curve. The second component velocities are $0.1$, $0.05$, and $0.15~{}c$, for the blue, green, and red curves, respectively. The old heating-rate prescription (Korobkin et al., 2012) is shown as dashed curves while the solid curves show the updated prescription (Rosswog & Korobkin, 2024). The physics behind the two models is the same apart from the heating-rate prescription.
@@
The detection and interpretation of kilonovae light curves are heavily reliant on available kilonova models. Specifically, such models form a key aspect for designing surveys, understanding depth requirements for follow-up, and ultimately characterising and inferring the properties of candidate events. However, kilonova light curve models are subject to a number of different ingredients, each with its own underlying uncertainty. One of these critical modelling ingredients is the nuclear heating rate, i.e., the energy produced per second by one gram of r-process material, which is ultimately subject to the decay channels of r-process nuclei (see e.g., Cowan et al., 2021, for an extensive review on various aspects of rapid neutron capture nucleosynthesis). We note that the term “heating rate” is often used to describe the combination of the “naked” heating rate and a thermalization efficiency. In this work, we use heating rate to refer to the “naked” heating rate.
@@
The temporal decay of a kilonova lightcurve has significant diagnostic power and may provide clues to the ejecta composition. If a broad distribution of nuclei is available and all are decaying exponentially, one expects a power-law dependence of the heating rate as a function of time (Way & Wigner, 1948; Li & Paczyński, 1998; Metzger et al., 2010b; Hotokezaka et al., 2017). If, instead, only a few isotopes dominate the heating rate, one expects distinctive “bumps” in the lightcurve that can provide constraints on the half-lives of the decaying species (Grossman et al., 2014; Martin et al., 2015; Lippuner & Roberts, 2015; Rosswog et al., 2018; Wanajo, 2018; Wu et al., 2017; Kasliwal et al., 2022).
@@
The specific decay channels of radionuclides that occur and their decay products determine how much energy is deposited in the expanding ejecta cloud and the overall heating rate. Any neutrino products escape the ejecta and do not contribute to the overall heating rate, while the heating-rate efficiency of $\gamma-$rays decays very rapidly and becomes negligible after tens of days (Hotokezaka et al., 2016). On the other hand, charged particles, such as $\alpha$-particles and fission products, keep depositing energy into the ejecta even at late times, and this makes kilonova lightcurves sensitive to the contribution of those decay channels (Barnes et al., 2016; Rosswog et al., 2017; Kasen & Barnes, 2019; Wu et al., 2019; Zhu et al., 2018; Vassh et al., 2019; Giuliani et al., 2020).
@@
Understanding what decay channels take place and what their decay products are is non-trivial. To complicate matters further, the interactions themselves are sensitive to the overall composition of the kilonova ejecta, which itself is linked in a multitude of ways to the progenitor binary neutron star system, equation of state, and the fate of the remnant (e.g., Rosswog et al., 2018; Sarin & Lasky, 2021). For relatively high $Y_{e}$ values, $\beta$-decays dominate, and only the resulting electrons and photons are relevant for heating the ejecta. In low $Y_{e}$ ejecta, actinides can be produced which, via $\alpha$-decays and fission, can deposit large amounts of energy, altering the heating rate by up to an order of magnitude depending on the underlying nuclear mass model (Barnes et al., 2016; Rosswog et al., 2017; Wu et al., 2019; Zhu et al., 2021; Barnes et al., 2021; Lund et al., 2023).
@@
In principle, kilonova models should be calculated self-consistently, such that feedback from the energy released and how this affects the expansion of the ejecta and nucleosynthesis/heating rate is incorporated. Recent work suggests that failure to include this feedback can compromise the abundance evolution (Magistrelli et al., 2024). Moreover, kilonova models should keep track of the exact decay channels to determine both the “naked” heating rate (i.e., the raw output) and the thermalization efficiency (the fraction that is coupled to the ejecta). This is difficult in practice since such models would need to be coupled to nuclear reaction networks which is computationally prohobitive. Instead, it is common to construct separate expressions for the naked heating rate, $\dot{q}$, and the thermalization efficiency, $\epsilon_{\rm th}$. The effective heating rate is then $\dot{q}_{\rm eff}=\epsilon_{\rm th}\;\dot{q}$.
@@
$\dot{q}=2q_{0}\left(\frac{1}{2}-\frac{1}{\pi}{\rm arctan}\frac{t-t_{0}}{\sigma}\right)^{\alpha},$ (1)
with $q_{0}=2\times 10^{18}$ erg/(g s), $t_{0}=1.3$ s, $\sigma=0.11$ s, $\alpha=1.3$ (Korobkin et al., 2012), consistent with heating-rate estimates from Metzger et al. (2010a), and the thermalization efficiency, $\epsilon_{\rm th}$ computed following Barnes et al. (2016) or Wollaeger et al. (2018). This specific heating-rate prescription was obtained as a fit to the heating rate from a broad set of Newtonian+GW-backreaction hydrodynamics simulations with the stiff equation of state of Shen et al. (1998a, b) and a multi-flavour leakage scheme (Rosswog & Liebendörfer, 2003). This prescription has served as the basis for a large variety of kilonova models, including the majority of model fits to the lightcurve of AT2017gfo, GRB211211A, and GRB230307A (Villar et al., 2017; Nicholl et al., 2021; Rastinejad et al., 2022; Levan et al., 2024), and detailed radiative transfer simulations (Bulla, 2019).
Figure 2: $95\%$ credible interval from $5000$ simulated kilonovae with fixed ejecta masses and temperature floors and varying velocities and opacities for old prescription in blue and updated prescription in green.
@@
Full general-relativistic simulations with more sophisticated neutrino treatments, produce, as expected, a broader range of electron fractions (Perego et al., 2019; Shibata & Hotokezaka, 2019), which are not adequately captured by the previous prescription (Rosswog & Korobkin, 2024). The fit in the latter reference is based on a library of parametrized trajectories where velocity and electron fraction were systematically varied. The actual nuclear network calculations were performed with WinNet (Winteler, 2012; Winteler et al., 2012; Reichert et al., 2023) which is based on the BasNet network (Thielemann et al., 2011). The functional dependence of the fit is more complicated than the above prescription, with multiple additional terms and written as,
@@
where the coefficients no longer have fixed values, but depend on the velocity and electron fraction of the ejecta. We refer to the appendix of Rosswog & Korobkin (2024) for the explicit coefficient expressions [^1] We note that successfully implementing this prescription and interpolating to different values of $Y_{e}$ and ejecta velocity requires the coefficients to be cast in a more numerically stable form and the interpolant to be built on $\dot{q}$ directly instead of using the functional form with interpolated coefficient terms. We refer the interested reader to the Redback implementation on GitHub https://github.com/nikhil-sarin/redback. The role of this updated heating-rate prescription has been explored in detailed radiative transfer simulations of kilonovae (Bulla, 2023), or used for predictions for surveys like LSST using the Vera Rubin Observatory (Setzer et al., 2023). However, it was not explored in detail for semi-analytical models. Moreover, it is unclear what the consequences of using the previous heating-rate prescription in fitting light curves are and what the systematic uncertainty for inferences of kilonova ejecta parameters due to uncertainties in the heating rate truly is.
As discussed in Rosswog & Korobkin (2024), the updated prescription produces substantially different heating rates for $Y_{e}\gtrsim 0.3$ and at times ($t\gtrsim 1\,\mathrm{d}$). Therefore, for the majority of this Letter, we will focus more on the impact of the updated heating rate prescription of Rosswog & Korobkin (2024) in this regime. In particular, in Sec. 2, we briefly describe a semi-analytical kilonova model implemented in open-source software, Redback (Sarin et al., 2023), highlighting the differences in individual light curves and a population due to the heating-rate prescription alone. We follow this by showcasing the dangers of using the previous heating-rate prescription to fit realistic, simulated kilonovae observations in Sec. 3. We then discuss the implications of uncertainty in the heating-rate prescriptions and conclude in Secs. 4 and 5, respectively.
@@ 2 Modelling and impact of heating-rate
Early modelling of the photometry of AT2017gfo made it abundantly clear that one single ejecta composition was inadequate for explaining the observations, with a clear evolution from a bright blue to a red source (Cowperthwaite et al., 2017; Kasen et al., 2017; Metzger, 2019). This led to the development of multi-component models, where each distinct ejecta component is characterized by a single opacity (Kasen et al., 2017). For the specific case of AT2017gfo, a two-component model with one lanthanide-poor, blue component and another lanthanide-rich red component proved sufficient to provide adequate fits to the observations (Villar et al., 2017), with extensions to include the effects of other emission mechanisms such as shock-cooling (Arcavi et al., 2017), effects of the hypermassive neutron star wind (Metzger et al., 2018), or interaction with a jet (Kasliwal et al., 2017), which served to improve fits to the observed data.
Although these extensions are relevant, the basic ingredients of the majority of semi-analytical models for kilonovae are broadly similar. We first have the ‘naked’ heating-rate prescription, which, for the majority of models follows Eq. 1. Only some fraction of this input energy is available to thermalize this ejecta, which is often taken from Barnes et al. (2016),
@@
Here, a, b, c are functions of the ejecta velocity and mass, with the overall function built as a fit to simulations (Barnes et al., 2016). The overall bolometric luminosity is then estimated following the Arnett model for supernovae (Arnett, 1982), with a temperature and wavelength independent gray opacity, which is estimated from the electron fraction, $Y_{e}$ following e.g., Tanaka et al. (2020). We note that this conversion is also a source of systematic uncertainty, which we will return to in later sections. As most comparisons to data are done through observations in specific filters, an assumption for the spectral energy distribution is also required, with the typical choice being a blackbody with a floor temperature such that the photospheric radius starts receding and the temperature remains constant once the floor temperature is reached. We note that the opacity and therefore $Y_{e}$, is often fixed to a particular value in several analyses previously performed in the literature (Cowperthwaite et al., 2017; Villar et al., 2017; Nicholl et al., 2021).
@@
The focus of this work is to highlight the impact of the ‘naked’ heating-rate prescription on kilonova lightcurves and show the impact of using an incorrect prescription and the systematic uncertainty due to our current uncertain knowledge of the true heating rate. Therefore, we limit our modelling to the standard approach as outlined above forgoing more detailed treatments of e.g., the thermalization efficiency and its dependence on the electron fraction (Shenhar et al., 2024; Guttman et al., 2024), which are critical to accurately capture kilonova lightcurves and infer ejecta properties from observations.
As we described in the introduction, the heating-rate prescription used by these models is based on simulations where the ejecta was predominantly at an electron fraction $Y_{e}=0.04$, i.e., (following the conversion from Tanaka et al. (2020, see their Table 1)), a grey opacity of $38\,\mathrm{g^{-1}~{}cm^{2}}$, in stark contrast to the opacities fixed for different components when fitting AT2017gfo, and other kilonova candidates. For example, the two-component fit for AT2017gfo in Villar et al. (2017) had gray opacities of $0.5\,\mathrm{g^{-1}~{}cm^{2}}$ and $10\,\mathrm{g^{-1}~{}cm^{2}}$. We note that as the heating rates are broadly consistent for $Y_{e}\lesssim 0.015$ due to fission cycling, the impact of the opacity is likely small for grey opacities $\gtrsim 30$. A different heating-rate prescription has a dramatic impact on the inferred properties of the ejecta (we explore this in more detail in Sec. 3). However, first we investigate the impact of the updated heating-rate prescription on the kilonova lightcurves themselves. Throughout this work, we assume the electron fraction to opacity relationship from Tanaka et al. (2020).
@@
In Fig. 1, we show three representative two component kilonova model magnitudes at $40\,\mathrm{Mpc}$. Here the difference is only due to the heating-rate prescription, with the old prescription (Korobkin et al., 2012) shown as dashed curves while the solid curve is the updated prescription (Rosswog & Korobkin, 2024). For all three models we fix the ejecta masses of each component to $0.02$ and $0.04~{}M_{\odot}$, and floor temperatures to $1500$ and $500$ K, respectively to isolate the impact of the heating rate. The blue and green curves show two component ejecta with electron fractions of $0.35$, $0.05$, and $0.25$ and $0.1$, while the red curves show both components for an electron fraction of $0.4$. For the ejecta velocities, we fix the first component velocity to $0.05$c for the blue and red curves, and $0.25~{}c$ for the green curve, while the second component velocities are $0.1$c, $0.05$c, and $0.15$c, for the blue, green, and red curves, respectively. The red curves demonstrate the significant impact of the heating-rate prescription, not only in changing the magnitude at peak by up to $1$ mag, but also the temporal evolution of the lightcurve, in particular the decline rate. In general, across a broad range of parameters, we find that the previous prescription produces brighter kilonovae at peak, but the late-time evolution is generally faster, although the parameters for the green curve indicate the new prescription is brighter.
@@
The above detailed lightcurves highlight specific cases where the updated heating-rate prescription leads to a markedly different lightcurve. However, it is also instructive, to look at the trend from a broader set of lightcurves. To understand the impact across a broad sample, we create $5000$ lightcurves drawn from uniform priors on the electron fractions between $0.05$ and $0.5$. Truncated normal priors for velocity with the first component with $\mu=0.15$c, $\sigma=0.1$c, and the second component with $\mu=0.1$c, $\sigma=0.05$c. For both components we fix the minimum and maximum velocities to $0.05$c and $0.35$c, respectively. We also fix the ejecta masses and temperature floors to $0.01$, $0.03~{}M_{\odot}$ and $3500$ and $1500$ K, again to isolate the impact of the heating-rate prescription. In reality, the properties of the ejecta, such as the masses, opacities, velocities are all correlated and link to the properties of the binary, e.g., through relations calibrated to numerical relativity simulations (Coughlin et al., 2019; Setzer et al., 2023). However, these relations are largely uncertain and can at times make nonsensical predictions outside the regime of the small subset of numerical simulations they are based on (Henkel et al., 2023), something we wish to avoid.
@@
In Fig. 2, we show the $95\%$ credible interval from the sample of $5000$ lightcurves, with the old prescription in blue and the updated heating-rate prescription in green. In other words, across our simulated set of lightcurves, at any given time the magnitude of the kilonova at $40\,\mathrm{Mpc}$ lies within the shaded band for each model at $95\%$ confidence. These lightcurve distributions highlight the broad impact of the updated heating-rate prescription, with the new prescription in general result in a dimmer kilonova at peak but with greater variation in decline rates. Another feature these lightcurve distributions highlight is the significantly tighter shape of the previous prescription (blue) i.e., the evolution of the lightcurve with the previous prescription is largely similar regardless of the opacity or ejecta velocity. This is in stark contrast to the updated prescription, where the lightcurve evolution depends sensitively on the opacity and ejecta velocity, a direct consequence of their impact on the input heating-rate.